OpenStudy (anonymous):

Holly missed 8 of 183 school days this year. Approximately what percent of school days was Holly in attendance? a. 4% b. 23% c. 96% d. 77%

OpenStudy (anonymous):

@arabpride

OpenStudy (anonymous):

A

OpenStudy (anonymous):

8/183=0.04=8%

OpenStudy (anonymous):

no its not

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

8/183=0.04=4%

OpenStudy (anonymous):

srry XD

OpenStudy (anonymous):

it was c

OpenStudy (anonymous):

>_>

OpenStudy (anonymous):

First, determine how many school days Holly attended. 183 - 8 = 175 Next, find the percentage of school days she attended. 175 ÷ 183 = 0.9562841530 96% Holly attended approximately 96% of school days.

OpenStudy (anonymous):

kk i didn't read it correctly me no like lectures

OpenStudy (arabpride):

Please close this post if you're done. :3 Thanks!

jimthompson5910 (jim_thompson5910):

Icedragon50 is on the right track. Icedragon50 has shown she missed 4% of the entire year, so she attended class 100% - 4% = 96% of the time. That's one way to get the answer.